An aluminium plate foxed in a horizontal position has a hole of diameter 2.000 cm. A steel sphere of diameter 2.005 cm rests on this hole. All the lengths refer to a temperature of 10 degree celsius. The temperature of the entire system is slowly increased. At what temperature will the ball fall down? Coefficient of aluminium is 23x10^-6/degree celsius and that of steel is 11x10^-6/degree celsius.

Grade:

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Cathy,

It will pass through the hole when the diameter of both of them are same.

Let  it happens at temperature t degree.

For steel, V=Vi(1+Υt)

Vi=initial vol

Y=coeff of vol expansion.

For Al A=Ai(1+βt)

Ai=initial area.

we have considered the vol of steel ball but the area of the Al plate as it's a 2D body.

again at the required temperature radii will be the same. hence putting v=4/3∏(r)^3 & A=∏(r)^2 we get

t=218.8 degree

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Cathy !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur